# Global analysis of fragmentation functions to charged hadrons with high-precision data from the LHC

### v1

@@ I Introduction
Thanks to the QCD factorization theorem, in the presence of hard scales, cross sections can be approximated as a convolution of perturbatively calculable short-distance scattering involving partons with universal long-distance functions Collins et al. (1989), including PDFs and FFs. Therefore, different processes involving FFs can be related, such as single inclusive hadron production in $e^{+}e^{-}$ annihilation (SIA), semi-inclusive deep-inelastic scattering (SIDIS), and proton-proton collisions. Tremendous efforts have been devoted to perturbatively calculable parts of these cross sections. The hard coefficient functions for SIA have been calculated at next-to-next-to-leading order (NNLO) in QCD in Refs. Rijken and van Neerven (1996, 1997); Mitov and Moch (2006); Blumlein and Ravindran (2006). For SIDIS, the next-to-leading order (NLO) corrections to the coefficients are given in Refs. Altarelli et al. (1979); Nason and Webber (1994); Furmanski and Petronzio (1982); Graudenz (1994); de Florian et al. (1998); de Florian and Rotstein Habarnau (2013). The NNLO corrections have been calculated recently in Refs. Goyal et al. (2023); Bonino et al. (2024). Approximate NNLO and N3LO corrections have also been obtained from expansion of the threshold resummation expressions Abele et al. (2021, 2022). The coefficient functions for single-inclusive hadron production at hadron-hadron collisions are known at NLO Aversa et al. (1989); de Florian (2003); Jager et al. (2003). In addition to single-inclusive hadron production, jet fragmentation at hadron colliders provides a direct means to probe the FFs as functions of momentum fraction. Numerical study of jet fragmentation at NLO has been carried out in Ref. Arleo et al. (2014). Analytic results for inclusive jet fragmentation have been calculated in both perturbative QCD with narrow jet approximation Kaufmann et al. (2015), and in soft-collinear effective theory (SCET) Chien et al. (2016). There also exists experimental measurements of hadron production in association with an isolated photon or $Z$ boson. The factorization formula for $Z+h$ production within SCET has been derived in Ref. Kang et al. (2019). Despite of non-perturbative nature of FFs, their scale dependence follows the Dokshitzer-Gribov-Lipatov-Altarelli-Parisi (DGLAP) evolution equation with time-like splitting kernels, which can be perturbatively calculated. These time-like splitting kernels have been calculated at ${\cal O}(\alpha_{s}^{3})$ in the strong coupling constant in Refs. Mitov et al. (2006); Moch and Vogt (2008); Almasy et al. (2012); Chen et al. (2021); Ebert et al. (2021); Luo et al. (2021).
@@
Equipped with perturbative QCD calculations and various experimental measurements, a comprehensive fit involving various data samples to extract FFs is feasible. The representative efforts can be found in the works of DSS de Florian et al. (2007), HKNS Hirai et al. (2007), AKK Albino et al. (2008), NNFF Bertone et al. (2018), MAPFF Khalek et al. (2021), and JAM Moffat et al. (2021). They are carried out at NLO in QCD with different data samples and different theoretical prescriptions. The HKNS analysis only includes SIA data, while AKK and NNFF analyses incorporate both SIA and proton-proton collision data. The MAPFF study uses SIA and SIDIS data. The DSS study uses data from SIA, SIDIS, and proton-proton collisions. For JAM collaboration, they performed a simultaneous fit of PDFs and FFs by including DIS, SIDIS, SIA, and Drell-Yan data. It is interesting to note that the AKK group has extended their studies to include hadrons of $K_{S}^{0}$, Λ($\Lambda(\bar{\Lambda})$, in addition to pion, kaon and (anti-)proton. There also exists determinations of FFs at NNLO with SIA data only Bertone et al. (2017); Soleymaninia et al. (2018), and at approximate NNLO with SIA and SIDIS data Borsa et al. (2022a); Abdul Khalek et al. (2022). Nevertheless, one of the least explored directions is the simultaneous fit by including different identified final states from different scattering processes Soleymaninia et al. (2021, 2019).
@@ II.2.1 Hadron collisions
These jet fragmentation datasets cover a wide kinematic region, as shown in Fig. 2, where only data points with LO hadron momentum fraction $z>0.01$ are displayed. The reference momentum $p_{T,{\rm ref}}$ of the initiating parton at LO in QCD is chosen as the transverse momentum of the photon/$Z$ boson for $\gamma/Z$-tagged hadron production. For ATLAS inclusive jets or dijet events, $p_{T,\rm{ref}}$ is the transverse momentum of the jet. The green dashed lines correspond to the cut $p_{T,h}>4$ GeV at LO in QCD, as mentioned in Sec. II.1.2, to ensure the validity of the leading twist factorization formalism and convergence of perturbative calculations.
@@ II.2.2 SIDIS
COMPASS06Adolph et al. (2017a, b) 17.3 540 ${\rm pb}^{-1}$ $x\in[0.14,0.4],y\in[0.3,0,5]$ $\pi,K,h$ $\frac{dM^{h}}{dz}$ 124
@@ V.1 Overall agreement
Figure 10: Histogram of the effective Gaussian variable $S_{E}$ for a total of 138 subsets of data and subsets of SIA/hadron in jet/$pp$ inclusive/SIDIS experiments. The blue and red curves represent a normal distribution and a Gaussian distributions with mean and standard deviation calculated from the ensemble of $S_{E}$.
@@ VII Discussion and Conclusions
We demonstrate the broad applications of the NPC23 FFs when combined with the FMNLO program. For instance, we calculate the average jet charges for QCD inclusive dijet production at the LHC and find good agreement with the experimental measurements. Additionally, we present predictions on $Z$-tagged jet fragmentation in $PbPb$ collisions at the LHC while neglecting final state medium effects, with more information given in appendix C. Meanwhile, we have published version 2.0 of the FMNLO program  [^2] https://fmnlo.sjtu.edu.cn/f̃mnlo/data/FMNLOv2.0.tar.gz, which now includes a module for calculations of hadron production in SIDIS at next-to-leading order in QCD. Details of the module and benchmark comparisons to existing results are also provided in appendix D. All theoretical predictions used in this analysis, including those for hadron collisions, SIDIS or SIA, can be reproduced with this program. Upon the completion of NPC23 for charged hadrons, an immediate follow-up study would be a global analysis of FFs to neutral hadrons. In a longer term, we are working on implementation of hard coefficient functions at next-to-next-to-leading order in QCD in the FMNLO program, especially given recent progresses on calculations of SIDIS coefficient functions Goyal et al. (2023); Bonino et al. (2024). This will pave the way towards a global analysis of FFs at NNLO in QCD.
Acknowledgments. The work of JG is supported by the National Natural Science Foundation of China (NSFC) under Grant No. 12275173. HX is supported by the NSFC under Grants Nos. 12035007, 12022512 by the Guangdong Major Project of Basic and Applied Basic Research Nos. 2020B0301030008, 2022A1515010683. YZ is supported by the NSFC under Grant No. U2032105. XS is supported by the Helmholtz-OCPC Postdoctoral Exchange Program under Grant No. ZD2022004.
@@ Appendix A Comparison to other groups
In this section, we present a comparative analysis of our fit results with those fragmentation functions from the DSS and NNFF groups. The comparisons of NLO fragmentation functions at 5 GeV summed over charges are depicted in Fig. 32. The DSS fits, sourced from DSS21 Borsa et al. (2022b), DSS17 de Florian et al. (2017), and DSS07 de Florian et al. (2007), provide fragmentation functions for $\pi^{\pm},K^{\pm}$ and $p/\bar{p}$. Note that DSS only publishes FFs with momentum fraction $z>0.05$. Correspondingly, the NNFF sets used are from NNFF1.0 Bertone et al. (2017). Uncertainties of FFs are not publicly available for DSS sets, whereas the error band of NPC23 is estimated using the Hessian method at a 68% confidence level. For NNFF sets, the uncertainties are estimated by the Monte Carlo method with FF replicas. The subfigures are organized by flavors of hadrons and quarks. Different rows correspond to FFs for $\pi^{\pm},K^{\pm}$ and $p/\bar{p}$, respectively. The first column displays FFs from constituent quarks, the second from other light quarks and the gluon, and the third from heavy quarks. As evident from Fig. 32, the uncertainties of NPC23 are notably smaller than those from NNFF due to the inclusion of a large variety of LHC data as well as the SIDIS data, which imposes significant constraints on the FFs.
@@ A.1 Fragmentation to $\pi^{\pm}$
For fragmentation of pions, NNFF assumes isospin symmetry, resulting in coinciding results for $\pi^{\pm}$ from $u$ and $d$ quarks. Both DSS and our fit allow for different normalization of the two constituent quarks, and our results show a moderate violation of isospin symmetry, though within uncertainties. Overall, our results agree well with those of DSS for constituent quarks to pions FFs but differ with those of NNFF, especially in large $z$ regions, where the differences can be larger than the uncertainties. For $s$ quark fragmentation, FFs from NNFF and DSS are close, while NPC23 is much larger, with a discrepancy of about 50$\sim$ 60%. In small $z$ regions, our result and NNFF show different trends, with NPC23 showing an increasing-then-decreasing pattern and NNFF displaying a continuously decreasing trend. For gluon fragmentation, FFs differ significantly in small $z$ regions, especially where NNFF shows large oscillations with a substantial uncertainty that even dips below zero. However, at $z>0.3$, NNFF and NPC23 agree well, while DSS exhibits larger values. For heavy quark fragmentations ($c,b$), FFs from different fits agree well for $z>0.1$ and differ notably at small $z$ regions, accompanied by large uncertainties.
@@ A.2 Fragmentation to $K^{\pm}$
Fragmentation to $K^{\pm}$ from the $u$ quark is similarly described by fits from NPC23 and DSS, while NNFF displays a large discrepancy. Note NNFF assumes symmetry between $D_{u}^{K^{\pm}}$ and $D_{s}^{K^{\pm}}$. Fragmentation from the $s$ quark slightly deviates, but NPC23 and DSS show a similar trend with a peak in the middle, and the peak locations are close. For fragmentation from the $d$ quark, NPC23 aligns well with DSS. NPC23 and NNFF are close in the small $z$ region ($z<0.05$), after which NNFF starts oscillating and even becomes negative. For gluon fragmentation, DSS and NPC23 agree at $z\sim 0.5$, with significant discrepancies in other regions, and NNFF consistently shows a large discrepancy. For fragmentation from $c$ quark, the three results are quite different, with DSS being the largest, NPC23 in between, and NNFF being the smallest. For fragmentation from $b$ quark, DSS and NPC23 are close and larger than NNFF.
@@ A.3 Fragmentation to $p/\bar{p}$
Figure 32: Comparison of our NLO fragmentation functions to those from NNFFs and DSS at $Q=5$ GeV. The DSS fits have a limit of 0.05 in the momentum fraction, with $\pi^{\pm}$, $K^{\pm}$ and $p/\bar{p}$ results from DSS21 Borsa et al. (2022b), DSS17 de Florian et al. (2017) and DSS07 de Florian et al. (2007), respectively. The NNFF sets used are from NNFF1.0 for pion, kaon and proton respectively. The estimated uncertainties of FFs are also shown for NNFFs and for our fit. For $\pi^{\pm}$ and $p/\bar{p}$ the left (middle) panel shows results for $u$ and $d$ ($s$ and $g$). For $K^{\pm}$ the left (middle) panel shows results for $u$ and $s$ ($d$ and $g$). The right panels show results for $c,b$ quarks.
Figure 33: Similar to Fig. 32 but for $\pi^{+},~{}K^{+},~{}p$ at 5 GeV
A.4 Fragmentation to $\pi^{+}$, $K^{+}$, $p$
In Fig. 33, a comparative analysis of fragmentation functions to positively charged hadrons is presented, similar to the arrangement depicted in Fig. 32. This allows for a clear distinction between fragmentation from favored and unfavored quarks. The NNFF exhibits even larger uncertainties in this case since they use SIA data alone, which cannot separate fragmentation from quarks and anti-quarks. We also observe that FFs to $h^{\pm}$ and $h^{+}$ are not fully correlated for NNFF, as should be obtained from charge conjugation. For instance, the profiles for $D_{g}^{K^{+}}$ and $D_{g}^{K^{\pm}}$ exhibit distinct trends. Most characteristics among different groups resemble those observed for FFs summed over charges. The agreement between the three groups is worse for fragmentation from unfavored quarks in general, for instance, as for the $\bar{u}$ quark shown in the middle column.

### v2

@@ I Introduction
Thanks to the QCD factorization theorem, in the presence of hard scales, cross sections can be approximated as a convolution of perturbatively calculable short-distance scattering involving partons with universal long-distance functions Collins et al. (1989), including PDFs and FFs. Therefore, different processes involving FFs can be related, such as single inclusive hadron production in $e^{+}e^{-}$ annihilation (SIA), semi-inclusive deep-inelastic scattering (SIDIS), and proton-proton collisions. Tremendous efforts have been devoted to perturbatively calculable parts of these cross sections. The hard coefficient functions for SIA have been calculated at next-to-next-to-leading order (NNLO) in QCD in Refs. Rijken and van Neerven (1996, 1997); Mitov and Moch (2006); Blumlein and Ravindran (2006); Gehrmann and Stagnitto (2022); Bonino et al. (2024a).
For SIDIS, the next-to-leading order (NLO) corrections to the coefficients are given in Refs. Altarelli et al. (1979); Nason and Webber (1994); Furmanski and Petronzio (1982); Graudenz (1994); de Florian et al. (1998); de Florian and Rotstein Habarnau (2013). The NNLO corrections have been calculated recently in Refs. Goyal et al. (2023); Bonino et al. (2024b). Approximate NNLO and N3LO corrections have also been obtained from expansion of the threshold resummation expressions Abele et al. (2021, 2022). The coefficient functions for single-inclusive hadron production at hadron-hadron collisions are known at NLO Aversa et al. (1989); de Florian (2003); Jager et al. (2003). In addition to single-inclusive hadron production, jet fragmentation at hadron colliders provides a direct means to probe the FFs as functions of momentum fraction. Numerical study of jet fragmentation at NLO has been carried out in Ref. Arleo et al. (2014) and recently in Refs. Liu et al. (2023); Zidi et al. (2024); Caletti et al. (2024) due to development on QCD subtraction method.
@@
Analytic results for inclusive jet fragmentation have been calculated in both perturbative QCD with narrow jet approximation Kaufmann et al. (2015), and in soft-collinear effective theory (SCET) Chien et al. (2016). There also exists experimental measurements of hadron production in association with an isolated photon or $Z$ boson. The factorization formula for $Z+h$ production within SCET has been derived in Ref. Kang et al. (2019). Despite of non-perturbative nature of FFs, their scale dependence follows the Dokshitzer-Gribov-Lipatov-Altarelli-Parisi (DGLAP) evolution equation with time-like splitting kernels, which can be perturbatively calculated. These time-like splitting kernels have been calculated at ${\cal O}(\alpha_{s}^{3})$ in the strong coupling constant in Refs. Mitov et al. (2006); Moch and Vogt (2008); Almasy et al. (2012); Chen et al. (2021); Ebert et al. (2021); Luo et al. (2021).
Equipped with perturbative QCD calculations and various experimental measurements, a comprehensive fit involving various data samples to extract FFs is feasible. The representative efforts can be found in the works of DSS de Florian et al. (2007, 2015, 2017); Borsa et al. (2022a), HKNS Hirai et al. (2007), AKK Albino et al. (2008), NNFF Bertone et al. (2018), MAPFF Khalek et al. (2021), and JAM Moffat et al. (2021).
@@
They are carried out at NLO in QCD with different data samples and different theoretical prescriptions. The HKNS analysis only includes SIA data, while AKK and NNFF analyses incorporate both SIA and proton-proton collision data. The MAPFF study uses SIA and SIDIS data. The DSS study uses data from SIA, SIDIS, and proton-proton collisions. For JAM collaboration, they performed a simultaneous fit of PDFs and FFs by including DIS, SIDIS, SIA, and Drell-Yan data. It is interesting to note that the AKK group has extended their studies to include hadrons of $K_{S}^{0}$, Λ($\Lambda(\bar{\Lambda})$, in addition to pion, kaon and (anti-)proton. There also exists determinations of FFs at NNLO with SIA data only Bertone et al. (2017); Soleymaninia et al. (2018), and at approximate NNLO with SIA and SIDIS data Borsa et al. (2022b); Abdul Khalek et al. (2022). Nevertheless, one of the least explored directions is the simultaneous fit by including different identified final states from different scattering processes Soleymaninia et al. (2021, 2019).
@@ II.2.1 Hadron collisions
These jet fragmentation datasets cover a wide kinematic region, as shown in Fig. 2, where only data points with LO hadron momentum fraction $z>0.01$ are displayed. The reference momentum $p_{T,{\rm ref}}$ of the initiating parton at LO in QCD is chosen as the transverse momentum of the photon/$Z$ boson for $\gamma/Z$-tagged hadron production. For ATLAS inclusive jets or dijet events, $p_{T,\rm{ref}}$ is the transverse momentum of the jet. The green dashed line corresponds to the cut $p_{T,h}>4$ GeV at LO in QCD, as mentioned in Sec. II.1.2, to ensure the validity of the leading twist factorization formalism and convergence of perturbative calculations.
@@ II.2.2 SIDIS
COMPASS06Adolph et al. (2017a, b) 17.3 540 ${\rm pb}^{-1}$ $x\in[0.14,0.4],y\in[0.3,0,5]$ $\pi,K,h^{\pm}$ $\frac{dM^{h}}{dz}$ 124
@@ V.1 Overall agreement
Figure 10: Histogram of the effective Gaussian variable $S_{E}$ for a total of 138 subsets of data and subsets of SIA/hadron in jet/$pp$ inclusive/SIDIS experiments. The green and red curves represent a normal distribution and a Gaussian distributions with mean and standard deviation calculated from the ensemble of $S_{E}$.
@@ VII Discussion and Conclusions
We demonstrate the broad applications of the NPC23 FFs when combined with the FMNLO program. For instance, we calculate the average jet charges for QCD inclusive dijet production at the LHC and find good agreement with the experimental measurements. Additionally, we present predictions on $Z$-tagged jet fragmentation in $PbPb$ collisions at the LHC while neglecting final state medium effects, with more information given in appendix C. Meanwhile, we have published version 2.0 of the FMNLO program  [^2] https://fmnlo.sjtu.edu.cn/f̃mnlo/data/FMNLOv2.0.tar.gz, which now includes a module for calculations of hadron production in SIDIS at next-to-leading order in QCD. Details of the module and benchmark comparisons to existing results are also provided in appendix D. All theoretical predictions used in this analysis, including those for hadron collisions, SIDIS or SIA, can be reproduced with this program. Upon the completion of NPC23 for charged hadrons, an immediate follow-up study would be a global analysis of FFs to neutral hadrons. In a longer term, we are working on implementation of hard coefficient functions at next-to-next-to-leading order in QCD in the FMNLO program, especially given recent progresses on calculations of SIDIS coefficient functions Goyal et al. (2023); Bonino et al. (2024b). This will pave the way towards a global analysis of FFs at NNLO in QCD.
Acknowledgments. We thank Ignacio Borsa for providing the latest DSS pion fragmentation function grids. The work of J. G. is supported by the National Natural Science Foundation of China (NSFC) under Grant No. 12275173 and open fund of Key Laboratory of Atomic and Subatomic Structure and Quantum Control (Ministry of Education). H. X. is supported by the NSFC under Grant No. 12475139 and No. 12035007, and by the Guangdong Major Project of Basic and Applied Basic Research Grant No. 2020B0301030008 and No. 2022A1515010683. Y. Z. is supported by the NSFC under Grant No. U2032105 and CAS Project for Young Scientists in Basic Research Grant No. YSBR-117. X. S. is supported by the Helmholtz- OCPC Postdoctoral Exchange Program under Grant No. ZD2022004.
@@ Appendix A Comparison to other groups
In this section, we present a comparative analysis of our fit results with those fragmentation functions from the DSS, NNFF, JAM and MAP groups. The comparisons of NLO fragmentation functions at 5 GeV summed over charges are depicted in Figs. 32-34. To illustrate the difference between fragmentation from favored and unfavored quarks, we also show the FFs from different groups for positively charged hadrons in Figs. 35 and 36.
The DSS fits, sourced from DSS21 Borsa et al. (2022a), DSS17 de Florian et al. (2017), DSS14de Florian et al. (2015),and DSS07 de Florian et al. (2007), provide fragmentation functions for $\pi^{\pm},K^{\pm}$ and $p/\bar{p}$. Note that DSS only publishes FFs with momentum fraction $z>0.05$. For the uncertainties, the DSS21 release includes Monte Carlo uncertainties, while the DSS17 and DSS14 release incorporate Hessian uncertainties. Correspondingly, the NNFF sets used are from NNFF1.0 Bertone et al. (2017). For NNFF sets, the uncertainties are estimated by the Monte Carlo method with FF replicas. For JAMMoffat et al. (2021) and MAPKhalek et al. (2021) FF sets, uncertainties are estimated with Monte Carlo method as well. The error band of NPC23 is estimated using the Hessian method at a 68% confidence level. As evident from Figs. 32-34, the uncertainties of NPC23 are notably smaller than those from NNFF due to the inclusion of a large variety of LHC data as well as the SIDIS data, which impose significant constraints on the FFs.
@@ A.1 Fragmentation to $\pi^{\pm}$
For fragmentation of pions, NNFF assumes isospin symmetry, resulting in coinciding results for $\pi^{\pm}$ from $u$ and $d$ quarks. Both DSS and our fit allow for different normalization of the two constituent quarks, and our results show a moderate violation of isospin symmetry, though within uncertainties. Overall, our results agree well with those of DSS for constituent quarks to pions FFs but differ with those of NNFF, especially in high-$z$ regions, where the differences can be larger than the uncertainties. The NNFF determination exhibits notably larger uncertainties compared to both NPC23 and DSS, particularly in the gluon and strange quark channels at low-$z$ region. Both DSS and NPC23 show relatively smaller and comparable uncertainty bands, suggesting more constrained determinations. For $s$ quark fragmentation, FFs from NNFF and DSS are close, while central value of NPC23 is much larger, with a discrepancy of about 50$\sim$ 60%. This difference becomes less significant when uncertainties are taken into account. In low-$z$ regions, our result and NNFF show different trends, with NPC23 showing an increasing-then-decreasing pattern and NNFF displaying a continuously decreasing trend. For gluon fragmentation, FFs differ significantly in low-$z$ regions, especially where NNFF shows large oscillations with a substantial uncertainty that even dips below zero. However, at $z>0.3$, NNFF and NPC23 agree well, while DSS exhibits larger values. For heavy quark fragmentations ($c,b$), FFs from different fits agree well for $z>0.1$ and differ notably at low- $z$ regions, accompanied by large uncertainties.
@@ A.2 Fragmentation to $K^{\pm}$
Fragmentation to $K^{\pm}$ from the $u$ quark is similarly described by fits from NPC23 and DSS, while NNFF displays a large discrepancy. The error bands of NPC23 and DSS overlap in intermediate-$z$ region ($0.70.1<z<$ 0.7). Note that NNFF assumes symmetry between $D_{u}^{K^{\pm}}$ and $D_{s}^{K^{\pm}}$. Fragmentation from the $s$ quark slightly deviates, but NPC23 and DSS show a similar trend with a peak in the middle, and the peak locations are close. For $u$ quark fragmentation to $K^{\pm}$, DSS and NPC23 analyses show agreement in most regions. In contrast, NNFF diverges significantly in intermediate-$z$ regions, displaying a much larger peak in the middle. For fragmentation from the $d$ quark, NPC23 aligns well with DSS and the error bands encompass DSS curve in most regions. NPC23 and NNFF are close in the low-$z$ region ($z<0.05$), after which NNFF starts oscillating and even becomes negative. For gluon fragmentation, DSS and NPC23 agree at $z\sim 0.5$, with significant discrepancies in other regions, and NNFF consistently shows a large discrepancy. For fragmentation from $c$ quark, the three results are quite different, with DSS being the largest, NPC23 in between, and NNFF being the smallest. But there are still slight overlaps between NPC23 and DSS in region $z<0.4$. For fragmentation from $b$ quark, results from three analyses are close and the error bands overlap in region $z>0.2$. But in low-$z$ region discrepancies persist.
@@ A.3 Fragmentation to $p/\bar{p}$
Figure 32: Comparison of our NLO fragmentation functions to pion with those from NNFF and DSS at $Q=5$ GeV. The DSS results are calculated by using DSS21 sets Borsa et al. (2022a). The NNFF results are from NNFF1.0 Bertone et al. (2017).
Figure 33: Same as Fig. 32 but for FFs to $K^{\pm}$. The DSS results are calculated by using DSS17 sets de Florian et al. (2017).
Figure 34: Same as Fig. 32 but for FFs to $p/\bar{p}$. The DSS results are calculated by using DSS07 sets de Florian et al. (2007).
Figure 35: Same as Fig. 32 but for FFs to $\pi^{+}$. The JAM20Moffat et al. (2021) and MAPFF10Khalek et al. (2021) sets are used respectively.
Figure 36: Same as Fig. 32 but for FFs to $K^{+}$. The JAM20Moffat et al. (2021) and MAPFF10Khalek et al. (2021) sets are used respectively.
A.4 Fragmentation to $\pi^{+}$, $K^{+}$
In Figs. 35 and 36 a comparative analysis of fragmentation functions to positively charged hadrons is presented. The fragmentation functions derived from NPC23, JAM Moffat et al. (2021), and MAP Khalek et al. (2021) analyses exhibit varying patterns of agreement and tension across different parton flavors for both $\pi^{+}$ and $K^{+}$ at $Q=5$ GeV. For favored quarks, JAM consistently predicts higher fragmentation probabilities compared to both NPC23 and MAP, with this feature being particularly pronounced in the pion sector. In contrast, the unfavored quark channels show better agreement among the three analyses, especially notable in the $\bar{u}\to K^{+}$ fragmentation. The gluon fragmentation functions demonstrate significant differences across all analyses for both mesons in terms of shape and magnitude, with MAP showing more pronounced fluctuations in the intermediate-$z$ region. For heavy quarks, all three analyses converge well within their uncertainties, particularly in the high-$z$ region, although some tension persists in the low-$z$ domain where JAM typically exhibits larger uncertainty bands. While the analyses show reasonable agreement in the high-$z$ region across most parton species, significant differences in both central values and uncertainty estimates remain in the low-$z$ region.